Main References

[A&B] C.D. Aliprantis and O. Burkinshaw, Problems in real analysis, Acad. Press, 1990.

[Aup] B. Aupetit, A primer on spectral theory, Springer-Verlag, 1991. [B&D] F.F. Bonsall and J. Duncan, Complete normed algebras, Springer-Verlag, 1973.

[Con] J.B. Conway, A course in functional analysis, New York, Springer-Verlag, 1990. [G&O] B.R. Gelbaum and J.M.H. Olmsted, Theorems & counterexamples in mathematics, Springer-Verlag, 1990.

[Gui] A. Guichardet, Special topics in topological algebras, New York, Gordon & Breach, 1968.

[Hel] A.Ya. Helemskii, The Homology of Banach and topological algebras, Kluwer Acad. Pub., 1989.

[Hal] P.R. Halmos, A Hilbert space problem book, Princeton, Van Nostrand, 1967. [K&R1] R.V. Kadilon and J.R. Ringrase, Fundamentals of the theory of operator algebras (I), Acad. Press, 1983. [K&R2] R.V. Kadilon and J.R. Ringrase, Fundamentals of the theory of operator algebras (II), Acad. Press, 1986. [K&R3] R.V. Kadilon and J.R. Ringrase, Fundamentals of the theory of operator algebras (III), Acad. Press, 1991.

[Ker] E. Kreyszig, Introductory functional analysis with applications, John Wiley & Sons, 1978.

[Mad] I.J. Maddox, Elements of functional analysis, Cambridge Unir.press, 1970.

[Mur] G.J. Murphy, C^{∗}-algebras and operator theory, Academic Press, 1990.

[Ped] G.K. Pederson, Analysis now, Springer-Verlag, 1989.

[Ric] C.E. Rickart, General theory of Banach algebras, Princeton, Van Nastrand, 1960.

[Rud1] W. Rudin, Functional analysis, McGraw-Hill, 1989.

[Rud2] W. Rudin, Real and complex analysis, McGraw-Hill, 1986.

[Sak1] S. Sakai, C^{∗}-algebras and W^{∗}-algebras, Springer-Verlag, 1971.

[Sak2] S. Sakai, Operator algebras in dynamical systems, Cambridge Univ. Press, 1991.

[Zel] W. Zelazko, Banach algebras, Elsevier Publishing Company, 1973.